OpenStudy (gmijares.01):

1. A 12 kg barrel is pulled up by a rope. The barrel accelerates at 1.2 m/s2. Find the force exerted by the rope. Show all your work. Im very confused

OpenStudy (angel.candy):

Total acceleration = (gravity + 1.2) = (9.8 + 1.2) = 11m/sec^2. Tension force = (mass x acceleration), = 12 x 11, = 132N.

OpenStudy (vipulrakta):

 in order to this problem t= tenion in the rope in +ve direaction of y axis a = acceleration also in +ve dreaction of y axis mg = weight of body _ve direaction of y axis F(net)= tension + weight \[F= T+(-mg)\] \[T=F+mg \] we know F= ma then tension \[T=m(g+a)\] put given value we get T=12(9.8+1.2) T=132